OpenStudy (anonymous):

What is the quotient of (2x2 + 14x + 24) ÷ (x + 4) ??

OpenStudy (anonymous):

2x^2 +14x+24 2(x^2+7x+12) 2(x+3)(x+4) ----------------- x+4 2(x+3)

OpenStudy (anonymous):

2x^2+14x+24= 2(x+3)(x+4) 2(x+3)(x+4)/(x+4)=2(x+3)

OpenStudy (anonymous):

you guys are the best!!!

OpenStudy (anonymous):

You're the best! I need some math practice over the summer.

OpenStudy (anonymous):

:D awe. thanks. im becoming your fan lol